News by Date
News by Category
Solar System
White Dwarfs
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Galaxy Clusters
Cosmology/Deep Field
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
Web Shortcuts
Chandra Blog
RSS Feed
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 19 February 2010 9.00am EDT

During the last week the observing schedule was replanned to include a 23 ks observation of the source U Scorpii that was accepted as a fast turn-around Target Of Opportunity (TOO) on Feb 12. Loads were replanned to remove observations of 0FGL J1821.4-1444, PGC1217593, and SDSS J144052.60-0235, which will be rescheduled in future weeks. The replanned loads were uplinked on Feb 13.

Chandra passed through a 35 minute, penumbra-only lunar eclipse on Feb 13 with nominal power and thermal performance. The eclipse was monitored in real-time. Real-time procedure were executed to disable SCS-29 and to set the short-eclipse flag to false (medium) on Feb 13 following the completion of the eclipse.

A Chandra press release was issued on Feb 17 describing new findings from Chandra that have provided a major advance in understanding a type of supernova critical for studying the dark energy that astronomers think pervades the universe. The results show mergers of two dense stellar remnants are the likely cause of many of the supernovae that have been used to measure the accelerated expansion of the universe. For details see:

The schedule of targets for the next week is shown below and includes an observation of VCS2 J0109+6133 accepted as a TOO on Feb 17 and observations of LHS 2065 and DE Tau coordinated with HST.

Radiation Belts Feb 21 VCS2J0109+6133 ACIS-S cl0422-5009 ACIS-S Feb 22 CygnusOB2 (3 obs) ACIS-I 4c+29.30 ACIS-S Feb 23 Radiation Belts Feb 24 CygnusOB2 (2 obs) ACIS-I 4c+29.30 ACIS-S Feb 25 M31 HRC-I Feb 26 LMCX-1 ACIS-S/HETG Radiation Belts LHS2065 ACIS-S Feb 27 NGC4472 ACIS-S HD119516 ACIS-S Feb 28 DETau ACIS-S


All spacecraft subsystems continued to support nominal operations.

Return to Status Reports